OpenStudy (anonymous):

the slope of the curve y^4=y^2-x^2 at the point (sqrt(3)/4 , 1/2)

OpenStudy (shamil98):

Differentiate implicitly then evaluate at x = sqrt(3/4) and y = 1/2

OpenStudy (dumbcow):

\[4y^3 y' = 2y y' + 2x\] \[y' = \frac{2x}{4y^3 - 2y}\] plug in point \[y' = \sqrt{3}\]